OpenStudy (peterg2001):

help will medal \ Add 13 5/12 +(-4 11/12) Reduce, if possible.

OpenStudy (peterg2001):

@dtuniyants

OpenStudy (anonymous):

write fraction 13 5/12 in standart form: 13+5/12=156/12+5/12=161/12 Reduce parathentles: 161/12-(4 11/12) Do the same thing with 4 11/12. 48/12+11/12=59/12. Now write the whole expression: 161/12-59/12=(161-59)/12=110/12